OpenStudy (fluffynose):

MEDALS MEDALS MEDALS MEDALS MEDALS! A circle is centered at the point (5, -4) and passes through the point (-3, 2). what the equation?

OpenStudy (faiqraees):

\[\large\rm (y-y1)^2+(x-x1)^2 =r^2 \] where (x1,y1) denotes the centre of the circle and r portrays the radius of the circle.

OpenStudy (fluffynose):

im not undering standing i got 100^2 but that translate into 10,000 but thats not and equation?

OpenStudy (fluffynose):

Can somebody help explain

OpenStudy (welshfella):

x1 and y1 are the center of the circle so replace x1 by 5 and y1 by -4 this gives you (x - 5)^2 + (y + 4)^2 = r^2 now you need to find the radius r

OpenStudy (faiqraees):

Now substitute the points it is passing with in place of x and y

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

this will give you the value of r^2 Then you are done.

OpenStudy (fluffynose):

okay so what i did was used the two points given for x and y. so it was like this when i wrote it out, (-4-2)^2+(5--3)^2= r^2 and i got 100^2 which is 10,000, im not quite understanding what to do next. did i mess up? could you do white board? i learn better like that

OpenStudy (welshfella):

r^2 comes to -8^2 + 6^2 = 64 + 36 = 100 so RHS =r^2 = 100

OpenStudy (welshfella):

sorry cant do white board

OpenStudy (fluffynose):

where'd you get -8

OpenStudy (welshfella):

(x - 5)^2 = (-3-5)^2 = (-8)^2

OpenStudy (welshfella):

- should have written it (-8)^2

OpenStudy (fluffynose):

Im not understanding did i do it right?

OpenStudy (welshfella):

correct is (-3 - 5)^2 + ( 2 + 4)^2 = (-8)^2 + 6^2 = 100

OpenStudy (welshfella):

you didnt substitute the values correctly It was pure luck you got it 100!

OpenStudy (welshfella):

the x is replaced by -3 and the y by 2

OpenStudy (fluffynose):

oh okay. haha Im sorry im diffucult i have a very hard time with mathh, now that i have 100, what exactly do i do with it?

OpenStudy (welshfella):

that is the value of r^2 in the general form of the equation so the final answer is (x - 5)^2 + ( y + 4)^2 = 100

OpenStudy (fluffynose):

i was definently over thinking lol Thank you! medals for you!

OpenStudy (welshfella):

yw

OpenStudy (fluffynose):

wait im on plato its fill in the blanks and it says this...

OpenStudy (fluffynose):

(x+___)^2+(y+___)^2=___

OpenStudy (fluffynose):

@welshfella

OpenStudy (welshfella):

well they are the values we just found for x1, y1 and r^2 5 , -4 and 100

OpenStudy (fluffynose):

okay thanks. :)